Please help me solve. Thank you
(1-x)^2=9
I think that I am supposed to set it to zero but it somehow isn't working for me.

Answer by checkley77(12844)

(1-x)^2=9
(1-x)=sqrt9
1-x=3
-x=3-1
-x=2
x=-2 ans.
Proof:
(1-(-2))^2=9
(1+2)^2=9
3^2=9
9=9